19 Almendarez Corporation is considering the purchase of a machine that would cost $160,000 and would...

19

Almendarez Corporation is considering the purchase of a machine that would cost $160,000 and would last for 4 years. At the end of 4 years, the machine would have a salvage value of $17,000. By reducing labor and other operating costs, the machine would provide annual cost savings of $42,000. The company requires a minimum pretax return of 7% on all investment projects. (Ignore income taxes.)

Click here to view Exhibit 12B-1 and Exhibit 12B-2, to determine the appropriate discount factor(s) using the tables provided.

The net present value of the proposed project is closest to: (Round your intermediate calculations and final answer to the nearest whole dollar amount.)
Answer #1
Present value of annual cash flows 142254 =42000*3.387
Present value of salvage value 12971 =17000*0.763
Total Present value 155225
Less: Investment cost 160000
Net present value -4775
Workings:
(1-(1.07)^-4)/0.07 = 3.387
1/1.07^4 = 0.763